# The Open Journal of Astrophysics Sulphur abundances in star-forming regions from optical emission lines: A new approach based on photoionization models consistent with the direct method

### v1

@@ 5 Summary and conclusions
The new version of HCm provides S/O ratios that are consistent with those derived using the direct method, with typical offsets below 0.05 dex and dispersions compatible with the observational uncertainties. As in previous versions, a good agreement between the abundances derived using the direct method and those calculated by HCm can be reached even in absence of any auroral line assuming an empirical relation between metallicity and the ionization parameter.
We demonstrated that the relation between O/H and U has a functional character that helps to derive abundances, but its physical meaning is much more related with a relation between metallicity and the hardness of the incident field of radiation, as already pointed out by Pérez-Montero et al. (2019b) using HCm-Teff.
@@
At low metallicities (12+log(O/H) 8.0), we observe greater scatter and a population of objects with elevated log(S/O), consistent with previous findings from IR-based analyses. These outliers may be linked to stochastic enrichment processes or variations in stellar yields.
@@ Acknowledgements.
This work has been funded by project Estallidos8 PID2022-136598NB-C32 (Spanish Ministerio de Ciencia e Innovacion). We also acknowledge financial support from the Severo Ochoa grant CEX2021-001131-S funded by MCIN/AEI/ 10.13039/501100011033. EPM also acknowledges the assistance from his guide dog Rocko without whose daily help this work would have been much more difficult. IAZ acknowledges funding from the Deutsche Forschungsgemeinschaft (DFG; German Research Foundation)—project-ID 550945879References

### v2

@@ 5 Summary and conclusions
The new version of HCm provides S/H abundances that are consistent with those derived using the direct method, with typical offsets below 0.05 dex and dispersions compatible with the observational uncertainties. As in previous versions, a good agreement between the abundances derived using the direct method and those calculated by HCm can be reached even in absence of any auroral line assuming an empirical relation between metallicity and the ionization parameter.
We demonstrated that the relation between O/H and U has a functional character that helps to derive abundances, but its physical meaning is much more related with a relation between metallicity and the hardness of the incident field of radiation, as already pointed out by Pérez-Montero et al. (2019b) using HCm-Teff. However, the assumption by HCm of an empirical anti-correlation between O/H and (i.e. the ionization parameter calculated with a fixed source) results useful to find out abundances without assuming multiple ionization sources in the models.
@@
At low metallicities (12+log(O/H) 8.0), we observe greater scatter and a population of objects with elevated log(S/O), consistent with previous findings from IR-based analyses. These outliers may be linked to stochastic enrichment processes or variations in stellar yields. However, the absence of these same points when S/O is compared with S/H, an indicator of total metallicity which is not affected by dust depletion, could indeed point out towards a large depletion factor in these objects.
@@ Acknowledgements.
This work has been funded by project Estallidos8 PID2022-136598NB-C32 and PID2022-136598NB-C33 (Spanish Ministerio de Ciencia e Innovacion). We also acknowledge financial support from the Severo Ochoa grant CEX2021-001131-S funded by MCIN/AEI/ 10.13039/501100011033. EPM also acknowledges the assistance from his guide dog Rocko without whose daily help this work would have been much more difficult. IAZ acknowledges funding from the Deutsche Forschungsgemeinschaft (DFG; German Research Foundation)—project-ID 550945879References